### v2

@@ 4.1 Pre-training improves performance on downstream endpoints up to a limit
Frist, we compare the model performances with varying pre-training dataset sizes. Figure 3 shows the performance of a model pre-trained with different increasing percentages of the GuacaMol [18] dataset molecules from 0% (i.e., randomly initialized wights) to 100%. The boxplots for the different pre-training subsets show that increasing size improves performance until 30% of the data as can be seen for solubility, permeability, HLM, and RLM. Significance analysis between the 0% and 30% setups supports these observations for solubility, permeability, and RLM (P-value $<0.05$, Suppl. Figure S2). However, models trained on further increments in pre-training data show similar and overlapping boxplots to the 30% setup (P-value >0.05)>0.05)> 0.05 ), Suppl. FigureS3) .
@@
Figure 3: The effect of increasing pre-training dataset size using DataSAIL splitting. 0% corresponds to a randomly initialized model with no pre-training and 100% correspond to the $\sim 1.3$M molecules of the GuacaMol dataset.
The hPPB and rPPB datasets pose a challenge to draw insights from due to the very low size of the datasets (i.e, $<200$ molecules), therefore, concluding a positive or negative result is not well supported.
Examining the values of the $R^{2}$ metric, one can observe that the permeability dataset showed high $R^{2}$ values as most of the values lie above 0.2, while most of the $R^{2}$ values for the rest of the datasets were below 0.15 (Figure 3). As a reminder, an $R^{2}$ value of 0 corresponds to a model as predictive as a model that predicts only the mean value, the upper bound is 1, which indicates a model significantly better than predicting the mean, and the lower bound is $\infty$- ∞, corresponding to a model worse than predicting the mean. Therefore, having most of the $R^{2}$ values at such low range imply that the different models still as naive as a mean-predictor despite the significance difference between the different setups.
Suppl. Figures S5 and S6 show comparable results under the more relaxed random splitting scheme. In these figures, the same trend holds as the performance improves significantly when pre-training on 30% of the pre-training data compared to a randomly initialized model (P-value $<0.05$ for all endpoints except plasma protein binding, Suppl. Figure S4), however, no significant improvement is observed beyond 30%. Finally, even though the random split is known to overestimate a model’s performance due to possible data leakage [55, 4], the $R^{2}$ values are still lying below 0.2 for solubility, HLM, and RLM, which indicates that the models are still as naive as a mean-predictor even for this relaxed split.
The lack of diversity can explain the limitedness of pre-training.
To better understand these findings, we explored the composition of the individual datasets in terms of molecular scaffold analysis. Table 3 shows the percentage of unique Murcko scaffolds in both pre-training (GuacaMol) and each endpoint datasets. The pre-training dataset consists of $\sim 1.3$M molecules and $\sim 407$K unique scaffolds. The percentage of unique scaffolds compared to the size of the dataset does not exceed 32%, which suggests that the pre-training dataset is low in diversity. This might explain the reason for the limited performance gain when using the full dataset. This is also inline with the findings from the neighboring field of material science where most databases were found to be redundant [29]. As a reminder, we select the different pre-training subsets by sampling from the different the dataset after clustering it based on tanimoto similarity and Morgan fingerprints using the BitBirch algorithm [32]. Therefore, most of the diversity of the dataset has been selected in the first 30% subset. Additionally, the table shows that the percentage of unique scaffolds per endpoint is extremely high (i.e., from 87% to 91%), while the percentage of common scaffolds with the pre-training dataset were between 31% and 36%. This suggests that the downstream molecules are too diverse on their own, and quite diverse from the pre-training dataset.
As a reminder, the objective used for pre-training is the masked language modeling (MLM), which exclusively learns patterns from the molecular structure as expressed by the tokenization of the SMILES sequence. Therefore, the low structural diversity of the pre-training dataset means that the model will learn frequent patterns quickly but it will struggle to generalize to structurally different sequences -as observed in our case-. This suggests that the MLM objective is not the most appropriate objective given the state of the datasets, which can also explain the low predictive power of the models.
% of Unique Murcko scaffolds ($\sim$) % of common scaffolds ($\sim$)
(GuacaMol $\cap$ endpoint)
GuacaMol 32
Solubility 87 34
Permeability 88 31
HLM 90 32
RLM 88 31
hPPB 91 32
rPPB 90 36
Table 3: Percentage unique Murcko scaffolds in pre-training and downstream datasets, as well as percentage of the common scaffolds between each endpoint moleculs and the pre-training dataset.
Conclusion:
Pre-training improved downstream endpoint predictions using 30% of the data. However, the usage of the full dataset did not provide significant improvements. This can be attributed to the low diversity of molecules — in form of Murcko scaffolds — in the pre-training dataset. Thus, adding more data does not necessarily introduce new patterns for the model to recognize. Generally, the model performances quantified by the $R^{2}$ metric were on the lower range for both the strict DataSAIL splitting and the relaxed random splitting, suggesting that the models are not much better than a simple model that only predicts the mean value. Further improvements in terms of pre-training dataset diversity and objective selection can help improve the models’ performance.
@@ 4.2 Domain adaptation improves performance significantly
Here, we compare domain adaptation with different objectives to the models without pre-training across the different pre-training dataset sizes. Figure 4 shows noticeable improvement with DA using the MTR objective for Solubility, Permeability, HLM, and RLM for all different pre-training setups. The CL objective also shows noticeable improvement compared to the No DA model, however, less prominent across the different pre-training setups compared to the MTR objective. Finally, DA using the MLM objective was not helpful as it performed comparable to the models without domain adaptation in all datasets. Significance analysis further corroborate these observations (Figure 5) as the DA models using MTR is showing P-values $<0.05$ for all pre-training setups for solubility, permeability, HLM and RLM, while DA using the CL objective shows P-values $<0.05$ for at least two of the different pre-training setups for the same datasets. DA using the MLM objective shows no significant difference to the No DA models for all datasets (P-value > 0.2).
Suppl. Figures S7 and S8 show comparable results for the random split experiments as DA using MTR and CL are both significantly better than the models with no domain adaptation for all datasets.
Figure 4: The effect of domain adaptation with different sizes of pre-training dataset using DataSAIL splitting. SE corresponds to the standard error of means. 0% corresponds to a randomly initialized model with no pre-training, and 100% corresponds to pre-training on the $\sim 1.3$M molecules of the GuacaMol dataset. No DA: a model with no domain adaptation, MTR: domain adaptation using multi-task regression of physicochemical properties, CL: domain adaptation using contrastive learning of the different representations of the SMILES string, and MLM: domain adaptation using masked language modeling.
Figure 5: Significance analysis of domain adaptation with different sizes of pre-training dataset using DataSAIL splitting as shown in Figure 4
Basic physicochemical properties might explain why MTR works.
DA analysis showed that MLM is not improving the model’s performance, CL is improving it to some extent, and MTR is improving it greatly. Here, we explain this behavior by touching back on the diversity feature of both datasets as well as the distribution of some basic physicochemical properties of the endpoint molecules. Table 3 shows that the molecules in the downstream enpoints are extremely diverse in terms unique Murcko scaffolds. Since MLM is an objectives that learns the sequence only, it will struggle to capture consistent patterns of the endpoint molecules, which can explain the weak performance of this objective during DA. The CL objective is also used as a structure-based objective like MLM, therefore, it is affected by the limitation in the structural composition of the dataset. However, the model sees multiple SMILES representations of the same molecule, which helps it capture patterns more consistently than MLM. This can explain the improved performance of the CL objective during DA compared to no DA at all or DA using the MLM objective. Suppl. Figures S12, S13, S14, S15, S16 and S17 illustrate the distribution of the molecules tested per endpoint for six physicochemical properties: molecular weight (MW), topological polar surface area (TPSA), the partition coefficient (LogP), number of hydrogen bond donors (HBD), hydrogen bond acceptors (HBA), and rotatable bonds. While the majority of molecules per endpoint possess unique Murcko scaffolds, they tend to cluster in a narrow range of histogram bins in the physicochemical property analysis (i.e., they share similar physicochemical profile). This explains why MTR is the best performing DA objective (Figures 4, S7, 5, and S8) as the molecular space is already easier to cluster by the physicochemical properties.
Conclusion:
Domain adaptation using the MTR objective significantly improved performance for three out of the four endpoints (solubility, permeability, and microsomal stability) compared to the models without domain adaptation. DA using the MLM objective did not yield significant improvements to the pre-trained-only models while the CL objective showed significant improvement in some setups. The MLM objective’s limited performance may be attributed to the high diversity in Murcko scaffolds for the downstream dataset. In contrast, the MTR objective was effective because the downstream data shared similar physicochemical profiles despite structural differences. Finally, the CL objective provided intermediate performance, improving upon the MLM objective by enabling the model to learn more consistent molecular representations.
4.3 MTR is better for pre-training, but best for domain adaptation.
Given the MTR objective’s promising performance during domain adaptation, we further explored if MTR would also outperform MLM during pre-training (Figure 6). The MTR objective consistently outperformed the MLM objective across all pre-training setups for solubility, permeability, HLM, and RLM. The same trend was observed with random splitting (Suppl. Figure S18). Additionally, the same trend of no significant improvement beyond 30% of the pre-training dataset was again observed when MTR was the pre-training objective (Suppl. Figures S19, S20, S21 and S22).
Figure 6: Performance comparison between the models pre-trained with MLM (blue) vs MTR (orange)objectives. DataSAIL splitting was used. 0% corresponds to a randomly initialized model with no pre-training, and 100% corresponds to pre-training on the $\sim 1.3$M molecules of the GuacaMol dataset.
Figure 7: comparing pre-training and domain adaptation objectives to identify the best approach. the models names follow the scheme of pre-training objective_pre-training dataset size_domain adaptation objective. DataSAIL splitting was performed.
Given that MTR is generally more expensive to optimize than MLM [22], and pre-training is inherently more computationally expensive than domain adaptation, it is crucial to allocate the best objective at the right stage to achieve optimal predictive power efficiently. Since pre-training with only 30% of the data consistently provides the best downstream performance, we focused on models with this setup. In the following, the performance of the models pre-trained with either MLM or MTR objective, and with either no domain adaptation or with MTR as a domain adaptation objective is investigated (Figure 7 and Suppl. Figure S23 for DataSAIL and random splitting, respectively). The figures show that models pre-trained with MLM and domain adapted with MTR (MLM_30%_MTR) perform similarly to those pre-trained with MTR, with or without domain adaptation. This is further supported by significance analysis (P-values $>$ 0.05, Suppl. Figures S25 and S24 for DataSAIL and random splitting, respectively). Given that MLM_30%_MTR is the most efficient setup without compromising performance, we conclude that MTR is best for domain adaptation rather than pre-training for the currently analyzed endpoints.
Basic physicochemical properties distributions of the upstream and downstream datasets can help understand this observation.
Our analysis demonstrated that the MTR objective for pre-training outperforms the MLM objective for predicting downstream properties. However, it did not improve performance when it was used in both pre-training and domain adaptation compared to it being used for domain adaptation only. To understand this behavior, we examined the distribution of basic physicochemical properties between the two datasets which showed that the downstream dataset spans similar distributions to the upstream dataset (Figure 8). Given that the two datasets share minimal molecules and exhibit significant diversity in Murcko scaffolds (Table 3), the MLM objective then proves insufficient for this scenario and the MTR objective is proven more helpful for pre-training. In both pre-training and domain adaptation, the upstream dataset spans much wider ranges than the downstream dataset (Figure 8), indicating that, although similar, the distributions still show many out-of-distribution molecules and physicochemical properties. Although the significance analysis indicates that none of the three models (MLM_30%_MTR, MTR_30%No DA, and MTR_30%_MTR) performs significantly better than the others, a visual inspection of the boxplots reveals that the distribution for the MLM_30%_MTR model is more compact, and consistently showing higher mean and median values. Thia observations leads us to suspect that the domain adaptation step for the model pre-trained with MTR is constrained by the patterns learned during pre-training. In contrast, domain adaption for models pre-trained with MLM appears to facilitate easier learning from the downstream dataset. This can help explain why MTR is more suited for domain adaptation rather than pre-training.
Figure 8: The distribution of basic physicochemical properties between the pre-training dataset (GuacaMol) and all the molecules in the downstream benchmark. The six physicochemical properties are molecular weight (MW), topological polar surface area (TPSA), the partition coefficient (LogP), number of hydrogen bond donors (HBD), hydrogen bond acceptors (HBA), and rotatable bonds.
Conclusion:
The results confirm that the MTR objective is more suitable than the MLM objective for the tested pre-training and downstream datasets. However, the added benefit of pre-training also stops after 30% of the dataset. Finally, although MTR is better as a pre-training objective, it is more efficient (without compromising performance) to use it during domain adaptation only and to employ the MLM objective during pre-training.
@@ Conclusion:
4.4 Our efficient transformer model rivals more complex models.
As examined so far, domain adaptation with the MTR objective using fewer samples ($\leq$ 3K molecules) is competitive against extensive pre-training beyond 400K molecules. The optimal model (i.e., MLM_30%_MTR) combines MLM for pre-training and MTR for domain adaptation, matching the predictive power of MTR for both stages. We compare the performance of the MLM_30%_MTR model to two other transformer models, MolBERT [1] and MolFormer [2]. MolBERT was pre-trained on $\sim$1.3M molecules using the GuacaMol dataset and three objectives were performed simultaneously during pre-training, MLM, MTR, and predicting whether two SMILES sequences correspond to the same molecule (SMILES-Eq). MolFormer was trained on 100M molecules compiled from the ZINC and PubChem databases while using only the MLM objective. A baseline is also shown for an RF model with default settings as provided by the scikit-learn library and trained on the physicochemical properties of the molecules.
The results in Figure 9 indicate that our model performs similar to MolBERT in all datasets, as evidenced by the close values of $R^{2}$ in the boxplots, and outperforms MolFormer for all datasets except rPPB. However, the significance analysis reveals that the three models perform similarly for all datasets (P-values $>0.05$), except for permeability, where both our model and MolBERT significantly outperform MolFormer (P-value = 0.048, Figure 10).
While our efficient model demonstrates competitive performance compared to more complex transformer models, both our model and MolBERT perform similarly to the simple RF + PhysChem model across all datasets. MolFormer also achieves similar results to the RF + PhysChem model for all datasets except permeability and RLM, where the simple RF model performs significantly better (P-value = 0.048).
Similar trends hold for all compared models when the random splitting scheme is used (Suppl. Figures S10 and S11)
Figure 9: Performance of the most efficient model from our analysis to models from the literature. MLM_30%_MTR corresponds to a transformer model pre-trained with 30% of the GuacaMol dataset using the MLM objective then domain adapted on the corresponding endpoint using the MTR objective. MolBERT [1] and MolFormer [2] are two transformer models from the literature. RF + PhysChem corresponds to a model trained on the physicochemical properties calculated by the RDKit package. The results are shown for the DataSAIL splitting scheme.
Figure 10: Significance analysis of the benchmarking shown in figure 9.
Yet opportunities for improvement remain
Our model that was pre-trained on fewer than $400$K molecules using the MLM objective and domain adapted on fewer than $3$K molecules with the MTR objective rivals the more complex and expensive MolBERT[1] and MolFormer[2] models. This suggests that current transformer models may be over-engineered and not optimized for the underlying chemical space. We already showed that the GuacaMol dataset is redundant, meaning it contains many very similar molecules, and is quite dissimilar to the endpoints molecules in terms of Murcko scaffolds, which could be limiting the benefit of pre-training. This affects both our models and MolBERT as both were pre-trained on the same dataset. MolFormer has been trained on 100M molecules combined from both ZINC and PubChem, therefore, spanning bigger and different chemical space than GuacaMol. All three transformer models have similar numbers of trainable parameters (i.e., $\sim 90$M), but are pre-trained on different magnitudes of molecules ($\sim 400$K for our model, $\sim 1.3$M for MolBERT, and $\sim 100$M for MolFormer). Scaling analysis of neural networks already suggests that there should be a fixed relationship between increasing data set size and the number of parameters, i.e. scaling laws, to achieve the desired improvement from such increases [56, 57]. Furthermore, a scaling experiment for molecular transformer models showed that molecular data exhibit a resolution-limited regime [58]. This regime suggests that a neural network trained on such data will need either a well-refined data set or a deeper neural network to capture enough nuances of the data [59]. Further analysis can investigate the best practices of model architecture and proper coverage of the chemical space.
Finally, a basic random forest (RF) model trained on physicochemical properties also showed competitive results to the three transformer models. We already showed that physicochemical properties are better descriptors of current downstream molecules than the structure alone. This is also highlighted when the same RF model was trained on the structural Morgan fingerprints, and it showed significantly worse performance than the RF model with physicochemical properties (Suppl. Figures S26, S27, S28 and S29). Our model, MolBERT and the RF models trained on the pyhsicochemical properties are outperforming models relying on structure-based representations like using Morgan fingerprints as input for RF or relying solely on MLM as in MolFormer. This shows that the main force behind the models’ performance is the molecular descriptors rather than the model architecture. Therefore, the current implementations of the transformer models is still limited in capturing fine-grained information about the molecules. Further investigations like probing and explainability can help identify the weak areas of the model’s learning, therefore, finding ways to further improve it.
Conclusion:
The results show that training a competitive transformer model can be done as efficiently as pre-training on $<400$K molecules with a simple objective like MLM, and domain adaptation on a couple of thousands of endpoint molecules using the MTR objective. Such a model can perform similarly to a highly domain-oriented model like MolBERT (i.e., using multiple domain objectives during pre-training) and better than a model pre-trained with a huge generic corpus like MolFormer (i.e., trained on 100M molecules from both ZINC and PubChem). However, the results also show that these three models, regardless of their complexity, are still on par with a basic RF model with default settings using physicochemical properties as molecular representation. We showed multiple reasons that can explain these observations, and we believe that the transformer models still have a large room for improvement to better represent molecular data and predict properties.

### v3

@@ 4.1 Pre-training improves performance on downstream endpoints up to a limit
Frist, we compare the model performances with varying pre-training dataset sizes, i.e, 0%, 30% ($\sim 400K$), 60% ($\sim 800K$), and 100% ($\sim 1.3M$) molecules.
Figures 3 and 4 show the MAE and $R^{2}$ performance of these models on seven independent molecular property datasets. The figures show that, for all datasets, at least two pre-training setups provided significantly better predictions than using a randomly initialized model (P-val < 0.01). However, the figures also show that, for all datasets, the 100% setup was either matched in performance (P-val > 0.05 for lipophilicity, HLM, and HPPB) or outperformed (P-val < 0.01 for permeability, the two solubility datasets, and RLM) by the 30% or the 60% setups. This shows that increasing pre-training dataset size plateaus, and further training is not recommended.
[…]
@@
Figure 3: MAE performance for increasing pre-training dataset sizes using Butina splitting. 0% corresponds to a randomly initialized model with no pre-training and 100% correspond to the $\sim 1.3$M molecules of the GuacaMol dataset. Two-tailed significance analyses were performed, therefore, the arrows in the heatmap helps recognizing the model with the improved performance. CI = confidence interval for the estimation of the mean.
Figure 4: $R^{2}$ performance for increasing pre-training dataset sizes using Butina splitting. 0% corresponds to a randomly initialized model with no pre-training and 100% correspond to the $\sim 1.3$M molecules of the GuacaMol dataset. Two-tailed significance analyses were performed, therefore, the arrows in the heatmap helps recognizing the model with the improved performance. CI = confidence interval for the estimation of the mean.
While significance testing assesses whether observed performance differences between models are statistically meaningful, it does not capture the practical relevance of these differences. A comprehensive evaluation therefore also requires examining the absolute values of the performance metrics. Mean Absolute Error (MAE), being unbounded and dataset-dependent, can be difficult to interpret without contextual knowledge. In such cases, experts in the specific application domain are typically better positioned to assess whether an MAE is acceptable.
Figure 4 shows the results for the $R^{2}$ metric, which is bounded within $[-\infty,1]$. Negative $R^{2}$ values indicate worse performance than simply predicting the mean, and a value of 1 corresponds to perfect prediction. For each dataset, the average $R^{2}$ in repeated runs remains notably low, for example, 0.08 for SolubilityAstrazeneca_ChEMBL, 0.06 for HLM, and as low as 0.02 for HPPB. The highest reported value is 0.33 for Permeability. Although statistical significance is observed, the absolute values of a bounded and interpretable metric highlight the remaining challenges in capturing the complex relationship between molecular structure and properties.
These findings highlight the critical need and opportunities for advancing transformer-based architectures in molecular property prediction through means other than dataset scaling alone.
Finally, the results presented in Figures 3 and 4 are based on train/test dataset splitting using Butina clustering with Morgan fingerprints, which is generally considered a stricter and more realistic evaluation strategy compared to random splitting. Figures S3 and S4 display the results from random splitting, which paint a similar overall picture to that observed with Butina clustering. While pre-training proves beneficial, performance tends to plateau around the 30% or 60% pre-training levels.
Furthermore, although the $R^{2}$ values under random splitting are generally higher than those observed with Butina clustering — as expected due to the less stringent data separation, they remain relatively low overall. The highest average $R^{2}$ value was observed for Lipophilicity, with an $R^{2}$ of 0.39.
Achieving comparable results under the more relaxed random splitting strategy further emphasizes the need for more effective approaches to model the structure–property relationship.
@@ 4.2 Domain adaptation improves performance significantly
Figure 5: MAE performance for a baseline model trained with pre-training only (No DA), and three models incorporating domain adaptation (DA) using different objectives: Masked Language Modeling (MLM), Contrastive Learning (CL), and Multi-task Regression (MTR) for physicochemical properties. P-values are from one-tailed paired t-tests comparing each DA model to the No DA baseline, under the hypothesis that DA improves performance. Significance levels: * $p<0.05$, ** $p<0.01$, *** $p<0.001$.
For the DA studies, we select the model trained on 30% of the Guacamol dataset as an efficient baseline and compare domain adaptation using three different objectives: masked language modeling (MLM), contrastive learning (CL), and physicochemical multi-task regression (MTR). Figure 5 presents the MAE performance for these four models. The results show that domain adaptation using the MTR objective consistently yields highly significant improvements across all datasets ($p$-value < $4e^{-9}$). The CL objective also demonstrates significant gains for all datasets except SolubilityAstrazeneca_ChEMBL. In contrast, the MLM objective outperforms the baseline model without domain adaptation on only two datasets, Lipophilicity and Permeability. Figure 6 presents the corresponding $R^{2}$ values, offering a diagnostic view of model performance. Notably, with domain adaptation using the MTR objective, the average $R^{2}$ values increased by around 0.2 units with the lowest average at $\sim 0.2$ , a substantial improvement over the near-zero values observed without adaptation. However, the maximum $R^{2}$ remains around 0.4, which is still relatively low and suggests that the model continues to face challenges in fully capturing the underlying variability in the data. These results demonstrate that chemically informed domain adaptation strategies, particularly those leveraging physicochemical properties, offer a more promising path for improving transformer-based molecular property prediction than scaling data alone.
The same trends are observed under the more relaxed random split, as shown in Figures S5 and S6. Domain adaptation using the MTR objective again yields consistent and highly significant performance improvements across all datasets. For the random-split scenario, the $R^{2}$ values increase from 0.2 to as high as 0.6.
Figure 6: $R^{2}$ performance for a baseline model trained with pre-training only (No DA), and three models incorporating domain adaptation (DA) using different objectives: Masked Language Modeling (MLM), Contrastive Learning (CL), and Multi-task Regression (MTR) for physicochemical properties. P-values are from one-tailed paired t-tests comparing each DA model to the No DA baseline, under the hypothesis that DA improves performance. Significance levels: * $p<0.05$, ** $p<0.01$, *** $p<0.001$.
4.3 MTR is a better pre-training objective, but best used for domain adaptation
Based on the strong performance of the MTR objective in domain adaptation, we further explored its use during pre-training and tested where it works best — pre-training only, domain adaptation only, or both. To this end, we pre-trained the same baseline model using the MTR objective alone, and subsequently applied MTR again for domain adaptation.
Figure 7 compares the performance of four models: one pre-trained with MLM only, one pretrained with MLM and adapted with MTR, one pre-trained with MTR only, and one pre-trained with MTR and adapted with MTR. The results show that pre-training with MTR consistently led to significantly better performance than MLM-based pre-training ($p$-value < 0.01). This highlights the added value of chemically informed objectives over generic, data-agnostic ones like MLM.
Interestingly, the best overall performance was observed for the model pre-trained with MLM and adapted with MTR (denoted as MLM_MTR). This model achieved significantly better results than the one pre-trained solely with MTR across all datasets except RLM, and also performed better than the model using MTR in both stages for all datasets except Permeability and RLM, where the performances were comparable.
These findings suggest that chemically informed objectives such as MTR are particularly effective when applied to data that is directly relevant to the downstream task. In this case, MTR yielded the greatest benefit when used during domain adaptation, where the model could leverage task-specific information. This underlines the importance of carefully aligning both the dataset and training objective with the prediction target in order to make the most of domain adaptation.
Figure 7: MAE performance to compare the best utilization of the MTR objective in either pre-training only, domain adaptation only, or both. MLM_No DA = Pre-training with MLM without domain adaptation. MLM_MTR = Pre-training with MLM and domain adaptation with MTR. MTR_No DA = Pre-training with MTR only. MTR_MTR = pre-training with MTR and domain adaptation with MTR. Two-tailed significance analysis were performed, therefore, the arrows in the heatmap helps recognizing the model with the improved performance. CI = confidence interval for the estimation of the mean.
@@ 4.3 MTR is a better pre-training objective, but best used for domain adaptation
4.4 Our efficient transformer model rivals more complex models, but not the simplest!
As examined so far, pre-training on fewer than 400K molecules combined with domain adaptation using the MTR objective on a relatively small number of samples ($\leq$ 4K molecules) yields better performance than extensive pre-training alone on more than 400K molecules. The optimal configuration, referred to as MLM_MTR, combines MLM for pre-training with MTR for domain adaptation.
We compare the performance of the MLM_MTR model to two recent transformer-based models: MolBERT [20] and MolFormer [26]. MolBERT was pre-trained on approximately 1.3 million molecules from the GuacaMol dataset, using a combination of three objectives: MLM, MTR, and SMILES equivalence (SMILES-Eq). MolFormer, on the other hand, was pre-trained on 100 million molecules from the ZINC and PubChem databases, using only the MLM objective. In addition, we include two baseline models based on Random Forest (RF) trained on either physicochemical descriptors or Morgan fingerprints.
The results shown in Figure 8 indicate that our model outperforms the Morgan fingerprint baseline across all datasets ($p$-value < 0.01), and also surpasses MolFormer on all datasets except HLM ($p$-value < 0.01). MolBERT achieves better performance than our model on only two datasets (Lipophilicity and Permeability), while performance is comparable on the remaining five datasets. These findings suggest that our lightweight model competes effectively with larger-scale models like MolFormer and MolBERT.
Interestingly, the strongest performer overall was the RF model trained on raw physicochemical descriptors, which achieved the best results on five datasets (Lipophilicity, Permeability, HLM, RLM, and HPPB) and performed comparably to MolBERT and our model on the remaining two solubility datasets. However, the $R^{2}$ values for the strongest model are still around 0.3 for most datasets (Solubility, HLM, RLM, and HPPB), and less than 0.6 for the remaining two datasets, lipophilicity and permeability (Figure S7.
This comparison reveals an important trend: the top-performing models in this study (RF + PhysChem, MolBERT, and MLM_MTR) incorporate explicit physicochemical information. In contrast, structure-based models (using Morgan fingerprints and MolFormer) showed comparatively lower performance. These results reinforce the importance of chemically and physically informed features for molecular property prediction, especially for ADME datasets. Moreover, the superior performance of the RF model on raw physicochemical descriptors, compared to transformer models using the same features as part of their objectives, suggests that these features were not fully exploited within the transformer architectures. This can potentially happen due to the architectural mismatch with low-dimensional, tabular data. While RFs are well-suited for this type of structured input, transformer models, originally designed for high-dimensional, unstructured data, may be over-parameterized and under-optimized for learning from a limited set of descriptors. This aligns with findings from recent studies showing that tree-based models often outperform deep learning models on tabular data without task-specific adaptations [58, 59].
Figure 8: MAE performance of the most efficient model from our analysis to models from the literature. MLM_MTR corresponds to a transformer model pre-trained with $\sim 400$K molecules using the MLM objective then domain adapted on the corresponding endpoint using the MTR objective. MolBERT [20] is a chemically aware transformer than has been pre-trained on $\sim 1.3$M molecules using MLM, MTR, and SMILES-EQ objectives. MolFormer [26] is a large-scale transformer pre-trained on 100M molecules using MLM. PhysChem and Morgan fingerprint correspond to two baselines using Random Forest models. Two-tailed significance analysis were performed, therefore, the arrows in the heatmap helps recognizing the model with the improved performance. CI = confidence interval for the estimation of the mean.